Financial Math

a). Altman Corporation has interest expenses of $120,000 annually. Altman’s annual sales are $4 million, its tax rate is 25%, and its net profit margin on sales is 10 percent. What is Altman’s TIE?

b). Back Alley Boys, Inc. had sales of $250,000, cost of goods sold of $80,000, depreciation expense of $27,000, and additions to retained earnings of $33,360. The firm paid out $30,000 in dividends. Assume a 34% income tax rate, what is the times interest earned ratio?

c). Using the following data, compute the value of cost of goods sold for Peterson Brewing:
Current liabilities=$340,000 quick ratio=1.8 inventory turnover=4.0 current ratio=3.3
Answer #1

SOLUTION :


a.


NPM = EAT / Sales = 10% 

=> EAT = Sales * 0.10 = 4000000 * 0.10 = 400000 ($) 

=> EAT = EBT ( 1 - 0.25) =  0.75 EBT

=> EBT = EAT / 0.75 = 400000 / 075 = 533333.33 ($)

=> EBIT = EBT + interest = 533333.33 + 120000 = 653333.33 ($)

=> TIE = EBIT / Interest = 653333.33 / 120000 = 5.44


So, Atman’s TIE is 5.44 (ANSWER).


b.


EAT = RE + Dividends = 33360 + 30000 = 63360 ($)


EBT = EAT / (1 - Tax rate) = 63360 / (1 - 0.34) = 96000 ($)


EBIT = Gross profit - Depreciation = (250000 - 80000) - 27000 = 143000 ($)


Interest = EBIT - EBT = 143000 - 96000 = 47000 ($)


So, TIE = EBIT / Interest = 143000 / 96000 = 1.49 


So,  Back Alley Boys’ times interest earned ratio (TIE) = 1.49 (ANSWER).


c.


Quick assets = CA - Inventory = Current Liabilities * quick ratio 

= 340000 * 1.8

= 612000 ($)


ITO = COGS / Inventory 

=> COGS = ITO * Inventory  = 4 * Inventory 


CR = CA / CL 

=> CA = CR * CL = 3.3 * 340000 = 1122000 ($)

=> Inventory = CA - Quick Assets = 1122000 - 612000 = 510000 ($)

=> COGS = 4 * Inventory = 4 * 510000 = 2040000 ($)


So, cost of goods sold (COGS) = 2040000 ($)   (ANSWER)

answered by: Tulsiram Garg